Two planes, which are 1580 miles apart, fly toward each other. Their speeds differ by 90mph. If they pass each other in 2 hours, what is the speed of each?
Answer by ikleyn(52799)

.

    2*x + 2*(x+90) = 1580   miles is the total distance.


(x is the speed of the slower plane; " 2 " represents 2 hours).


From the equation, step by step


    2x + 2x + 180 = 1580

    4x            = 1580 - 180 = 1400

     x = 1400/4 = 350 miles per hour is the speed of the slover,

                   and  350 + 90 = 440 miles per hour is the speed of the faster.

Solved.

-----------------

See introductory lessons on Travel and Distance problems
    - Travel and Distance problems
    - Travel and Distance problems for two bodies moving in opposite directions
    - Travel and Distance problems for two bodies moving in the same direction (catching up)
in this site.

They are written specially for you.

You will find the solutions of many similar problems there.

Read them and learn once and for all from these lessons on how to solve simple Travel and Distance problems.

Become an expert in this area.